Recent Log Entries

iocse4 restored

Log Number: 3858875Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-11-30 10:34:48 The output from the myRestore execution was as follows: iocse4 system@2020-11-30 10:34 2(2) restored seebpm@2020-11-30 10:34 2641(2641) restored SEE_BURT 61(61) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocse4 (success) Procedure return status: Success

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3858871Logbook(s): Body: 

Day #49, Week 11 Day 1, Plans:

Tunnel --
* Opportunistic cable pulls (CIS/EES-IC)

InjSB --
* Wien upgrade work, HV cabling, IN01B{00-03}
(Windham, Dela Cruz)
* CIS Rack work, IN01B{00-05} (Windham)
* Viewer cable terminations, IN03B07 (Norris)

Test Lab --
* Y-chamber prep for move into tunnel (Stutzman)
* Preparing for V-wien HV testing at UITF this week
(Hernandez-Garcia)
* Prebuncher RF frequency checks (Adderley, Wissman)
* A1/A2 cube assembly (Adderley)

Toolbox attended by: Hofler, Dela Cruz, Hernandez-
Garcia, Croke, Grames, Cole, Stutzman, Adderley,
Suleiman, Windham, Zhang

iocse4 restored

Log Number: 3858870Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-11-26 08:21:38 The output from the myRestore execution was as follows: iocse4 system@2020-11-26 08:21 2(2) restored seebpm@2020-11-26 08:21 2641(2641) restored SEE_BURT 61(61) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocse4 (success) Procedure return status: Success

Tsuneo's DAQ test plan

Log Number: 3858867Logbook(s): Entry Makers: poelkerBody: 

Tsuneo's plans to use the beams this Monday for the following purposes.

We've not run DAQ for more than 1 week due to the network traffic jam issues.
Would like to test how DAQ affects IBC network issues.

(1) Charge asymmetry (~40 Hz to DAQ) --- 3 hours or so (10 AM to 1 PM; time is just an idea)

Use keV beams to BPM and take scaler data to DAQ with the following conditions.

(a) Optical element IN (0 window delay) 1 micro A: 10 minutes

(b) Optical element IN (0 window delay) 0.5 micro A: 20 minutes

(c) Optical element IN (8 windows delay) 1 micro A: 10 minutes

Stop beams and get int

UITF locked up, today's plan

Log Number: 3858865Logbook(s): Entry Makers: poelkerBody: 

UITF locked up, gun ramping. Today we will check out the DAQ using an artifical charge asymmetry created using optical elements on the laser table, and then we will zero the charge asymmetry (i.e., remove the optical elements creating the charge asymmetry) and look for a physics asymmetry, beam on HD target. Tsuneo is the PI today.

3.5 hours into 125pA HDIce run

Log Number: 3858810Logbook(s): Entry Makers: poelkerBody: 

About 3.5 hours of beam at 125 pA, with a reduced holding field at IBC and modified raster amplitude. Very smooth running. Lock set to 105pA because some beam lost before dump, by design

30 minutes into new HDIce run

Log Number: 3858805Logbook(s): Entry Makers: poelkerBody: 

30 minutes into a new run 105pA, believed to be similar to 125pA run, now with new raster that results in some lost beam enroute to dump.

beam at HDIce now

Log Number: 3858799Logbook(s): Entry Makers: poelkerBody: 

Team HDIce is calculating new magnet holding field and raster settings, and then we do another 125pA run

buncher phase adjustment

Log Number: 3858764Logbook(s): Entry Makers: poelkerBody: 

At the last trip (BLM k401), I adjusted the buncher phase by four degrees. This helped improve the beam spot at ITVM904, minimize a tail. I don't think this adjustment will do anything for the ILMK401 BLM trips (we've had two today). I checked beam on keV viewers, thinkgs looked ok. 2 and 7 cell phases unchanged.

smooth operations at UITF: one hour into 125 pA run

Log Number: 3858758Logbook(s): Entry Makers: poelkerBody: 

One hour into the 125 pA HDIce target run. Smooth running.

In figure2, you see a jump in rad levels near A6, this happens when they perform an NMR measurement.

Yesterday, I believe Team HDIce learned the target performs better when colder. So I believe we are repeating the 125 pA run at a colder temp.

HDIce run at 125pA

Log Number: 3858757Logbook(s): Entry Makers: poelkerBody: 

Beam to HDIce at 125pA. Lock settings are shown. FE from gun about the same as before

locking up UITF now

Log Number: 3858751Logbook(s): Entry Makers: poelkerBody: 

Team HDIce is done transfering LHe from one dewar to another, locking up the cave now

Pages